FAR INFRARED FOURIER TRANSFORM SPECTROSCOPY IN EMISSION

O. PIRALI AND \underline{M. VERVLOET}, Laboratoire de Photophysique Moléculaire du CNRS, Bât. 350, Université Paris-Sud, 91405 Orsay Cedex, France.

We are currently developing sources allowing gas phase molecular emission to be recorded by FTS in the Far Infrared region. Our preliminary results were presented last year^{a}. Improvements of the efficiency of our sources result in the observation of lower frequency molecular emission. We illustrate this with two examples. By recording the thermal vibration-rotation emission spectra of few PAH’s using a new technique, the emission spectrum of anthracene was extended (for the first time to our knowledge) to its lowest frequency component located at around 87 cm^{−1}. By using a radio-frequency discharge through a flowing mixture of C_{2}H_{4} and N_{2}, rotational emission of HCN was detected in the spectral range 30 - 180 cm^{−1}. Rotational lines in pure bending motion were assigned up to the vibrational level (050). Results on HCN will be compared to those obtained by submillimeter-wave spectroscopy^{b} and infrared spectroscopy^{c}.

aO. Pirali and M. Vervloet, 58^{th} International Symposium on Molecular Spectroscopy, paper \underline{WE06} (2003).

bZelinger, T. Amano, V. Ahrens, S. Brünken, F. Lewen, H.S.P. Müller, and G. Winnewisser, J. Molec. Spectrosc., \underline{220}, 223 (2003).

cA. G. Maki, G. Ch. Mellau, S. Klee, M. Winnewisser, and W. Quapp, J. Molec. Spectrosc., \underline{202}, 67 (2000).